1. Project B has following cashflows at years 0 and 1 respectively: $ -4225 and $ 6302. What is the Internal Rate of Return (IRR) of this project?

2. Assume Fox Inc. has a P/E ratio of 22.5 and a stock price of $ 37.5. Calculate earnings per share of the company.
Answer #1

1)

IRR is the rate of return that makes initial investment equal to present value of cash inflows

Initial investment = FV / (1 + r)^n

4225 = 6302 / (1 + r)^1

(1 + r)^1 = 1.491598

1+ r = 1.491598

r = 0.4916 or 49.16%

IRR is 49.16%

2)

P/E ratio = Price / EPS

22.5 = 37.5 / EPS

EPS = 1.67